# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 3.3.5   Comments
If we count all observations from the previous discussion, there is just one missing resonance in the first and second excitation shells. However, five of these resonances have a $1^{*}$ or $2^{*}$ assignment only and their existence is not beyond doubt, and two are not (yet) included in the RPP (see Table 3.2). It is certainly mandatory that the existence of these states is solidified. The assignment of the observed states to multiplets is often a pure guess. Sometimes, the multiplets show a rather irregular pattern: For instance, the masses of the $\Delta^{*}$ quartet with $J^{P}=\nicefrac{{1}}{{2}}^{+},\cdots,\nicefrac{{7}}{{2}}^{+}$ are 1880$\pm$20, 1880$\pm$30, 1795$\pm$20, 1892$\pm$5 MeV, respectively. There is no obvious reason for the low mass of $\Delta(1905)\nicefrac{{5}}{{2}}^{+}$. The situation is even worse in the third excitation shell, since many resonances in the third shell are missing.
@@
The good agreement between the non-relativistic quark model predictions and the measured masses is, and has always been, extremely astonishing. Even constituent quarks should move at relativistic velocities when the nucleon is excited. Furthermore, the coupling of baryons to virtual baryon-meson intermediate states may lead to large baryon self-energies and to mixing of baryons of the same quantum numbers [552]. There is no reason at all to believe that the non-relativistic quark model may be useful in understanding the spectrum, except one: it works surprisingly well. Possibly, the concept of rapidly moving constituent quarks with a fixed rest mass is misleading: when a string evolves, a large fraction of the excitation energy may be stored in the string, the effective constituent quark masses increase, and the resulting motion could remain a classical one.
@@
In Ref. [553], two resonances were considered with one decay channel only, in which case one of the states becomes narrow and the other one broad. We know, however, no physical case corresponding to this situation. The second possibility are two close-by states with strong couplings to two decay modes. For example, the two mesons $K_{1A}$ and $K_{1B}$, the strange partners of $a_{1}(1260)$ and $b_{1}(1235)$, have the decay modes $K^{*}\pi$ and $K\rho$ which are expected with similar rates. Due to mixing, these two mesons repel each other, and the mixed state $K_{1}(1280)$ couples more strongly to $K\rho$ and $K_{1}(1400)$ very strongly to $K^{*}\pi$. On the other hand, these two scenarios – decays into one dominant channel or mutual repulsion of pairs of states – do not seem to be applicable in the high-lying baryon spectrum, where the number of negative-parity $N^{*}$ and $\Delta^{*}$ states expected in the third excitation shell is rather large (see Table 3.2). In the $N^{*}$ sector there is a low-mass doublet and in the $\Delta^{*}$ sector a low-mass triplet, and only these five low-mass resonances can be assigned to a unique multiplet. Above, there is exactly one observed state per partial wave from $J^{P}=\nicefrac{{1}}{{2}}^{-}$ to $\nicefrac{{9}}{{2}}^{-}$. In the quark model, the four $\Delta\nicefrac{{3}}{{2}}^{-}$ states above the $\Delta(1940)\nicefrac{{3}}{{2}}^{-}$ in the third excitation shell have calculated masses within a 100 MeV wide interval and the seven $N\nicefrac{{3}}{{2}}^{-}$ states above the $N(1875)\nicefrac{{3}}{{2}}^{-}$ fall into a 200 MeV wide interval, and these resonances are expected to have many decay modes.
@@
3.3.6   Hyperons
The simplest spectrum is that of the $\Omega$ family. All $\Omega^{*}$’s belong to the decuplet, which has fewer resonances than octet baryons; they all have charge $-1$, and they are narrower: the $\Omega(2012)$ with unknown $J^{P}$ (but most likely $\nicefrac{{1}}{{2}}^{-}$) has a width of less than 10 MeV, compared to 110 MeV for the $\Delta(1620)\nicefrac{{1}}{{2}}^{-}$. However, experimentally very little is known: Only four excited states are known, and two of them likely belong to higher bands.
Table 3.7 presents a quark-model interpretation of the $\Lambda$ and $\Sigma$ resonances [554] analogous to Fig. 3.2. In empty slots, no resonance is expected. Most exciting are the $\Lambda^{*}$’s, the only family that contributes to the flavor singlet spectrum. For each $N^{*}$ resonance, one $\Lambda^{*}$ octet state is expected. The $\Sigma^{*}$ resonances can belong to the $SU(3)_{f}$ octet or decuplet, so for each $N^{*}$ or $\Delta^{*}$ a $\Sigma^{*}$ resonance should exist. New data are expected from CLAS12 at JLab [555, 556] PANDA at GSI/FAIR [557], and ELSA [558]. For a recent review on hyperons we refer to Ref. [559].
Figure 3.7: (color online) Hyperon spectrum in the $SU(6)$ multiplet arrangement analogous to Fig. 3.2. In the first and second excitation band, all states expected from the quark model are shown, while the third band lists only multiplets for which at least one $\Lambda$ or $\Sigma$ candidate exists. The states with a dagger $\dagger$ are special: In the $\Lambda$ sector at about 2100 MeV, two states with $J^{P}=\nicefrac{{5}}{{2}}^{-}$ and $\nicefrac{{7}}{{2}}^{-}$ are expected with a mass separation of about 100 MeV but only one pair is known. In the $\Sigma$ sector, one pair is expected at about 1750 to 1800 MeV and two pairs at about 2000 to 2050 MeV. However, only two pairs are found and the third pair is missing. The observed pairs of states are possibly each mixtures of these configurations [554].
3.3.7   Baryon decays
Quark model calculations do not only give masses and widths of baryon resonances but also their wave functions (not only the harmonic oscillator basis wave functions). These can be used to calculate branching ratios, photocouplings or the $Q^{2}$ dependence of transition form factors. Table 3.5 gives a survey of references to calculations. A comprehensive review of the branching ratio calculations published before 2000 can be found in [560].
In baryon decays $N^{*}\to NM$, a new $q\bar{q}$ pair is created in a process as depicted in Fig. 3.8. The antiquark combines with a quark of the resonance to form a meson. The newly created $q\bar{q}$ pair does not form the meson $M$, which is forbidden by the OZI rule [561, 6, 562]. Often, the new $q\bar{q}$ pair is supposed to have the quantum numbers of the vacuum, $J^{PC}=0^{++}$. These ${}^{3}P_{0}$ models have met with considerable success.
Simple rules are at most approximately consistent with the experimental results. In hyperons, for example, the probability to emit a kaon is only 1/3, while the $s$ quark ends up in the baryon with a probability of 2/3. This makes the identification of hyperons in $K^{-}p$ elastic scattering difficult. In Table 3.6 we compare $N\bar{K}$ decays of $\Sigma$ hyperons with 4* rating with $N\pi$ decays of their non-strange partners. The suppression of $\Sigma^{*}\to N\bar{K}$ decays relative to their non-strange partners varies from 6 to 1.
In Refs. [563, 564], a novel attempt is made to gain insight into the internal structure of baryons by modeling real and imaginary parts of baryon resonances. For the fits, baryons are divided into several classes: $N$, $\Delta$, $\Lambda$, and $\Sigma$ on the one hand, and according to their naturality and signature [^2] The naturality is given by $\eta=+1$ if $P=(-1)^{J_{p}-1/2}$ and $\eta=-1$ if $P=-(-1)^{J_{p}-1/2}$; the signature $\tau$ by $\tau=\nicefrac{{P}}{{\eta}}$. on the other hand. The real parts of resonances in a given class show the expected linearity as functions of $J$, the imaginary parts a square root behavior. The square-root dependence is assigned to the contribution of the phase space to the scattering amplitude, which is proportional to the momentum $q\sim\sqrt{s-s_{t}}$. The hope was that major deviations from the observed pattern would signal important components beyond compact $3q$ physics. The authors of Ref. [563, 564] conclude that in the $\Lambda$ and $\Sigma$ sector a consistent picture emerges for the leading hyperon Regge trajectories. The $\Lambda(1405)$ will be discussed below (see Section 6.2), where its two-pole nature is discussed. In [563], it is found that the higher-mass pole of the $\Lambda(1405)$ fits into the general picture but the lower-mass pole does not, so it is suggested that it might be a molecule. The $N(1680)\nicefrac{{5}}{{2}}^{+}$, $N(1720)\nicefrac{{3}}{{2}}^{+}$, $\Delta(1930)\nicefrac{{3}}{{2}}^{-}$, and $\Delta(2400)\nicefrac{{9}}{{2}}^{-}$ are suggested to have sizable physics beyond the compact $3q$ picture. However, a few open questions remain: The $N$ and $\Delta$ are well-separated while the octet and decuplet $\Sigma$ resonances are not, and the decuplet $\Sigma(1385)\nicefrac{{3}}{{2}}^{+}$ is shown jointly with the octet $\Sigma(1775)\nicefrac{{5}}{{2}}^{-}$ (these quantum numbers are forbidden for a decuplet state) and with the decuplet $\Sigma(2030)\nicefrac{{7}}{{2}}^{+}$. The $N(1720)\nicefrac{{3}}{{2}}^{+}$ is related to the $N(1990)\nicefrac{{7}}{{2}}^{+}$; both belong to the second excitation shell, and both have $L=2$ as their leading orbital angular momentum. The $\Delta(1930)\nicefrac{{3}}{{2}}^{-}$ must have one unit of radial excitation, and the $\Delta(2400)\nicefrac{{9}}{{2}}^{-}$ likely has the same. It seems questionable whether these states should be related to the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$, and $\Delta(2420)\nicefrac{{11}}{{2}}^{+}$.
Table 3.5: Calculations of branching ratios BR, photocouplings $A$, and of the $Q^{2}$ dependence of transition form factors FF.
Ref. $A$ FF BR
[565, 566, 567] yes yes yes
[530, 568, 569] yes yes no
[570, 571, 572, 573, 574, 575, 576, 577, 578, 579] yes no yes
[580, 581, 582, 583, 584, 585, 586, 587, 588] no no yes
[589, 590, 591, 587, 592, 593] no yes no
$J^{P}$ $B^{*}\to N\pi$ $B^{*}\to N\bar{K}$ $B^{*}\to N\bar{K}$
$\frac{3}{2}^{-}$ $N(1520)$ (60$\pm$5)$\Sigma(1670)$ (9$\pm$3)$\Lambda(1690)$ (25$\pm$5)%
$\frac{5}{2}^{-}$ $N(1675)$ (40$\pm$2)$\Sigma(1775)$ (40$\pm$2)$\Lambda(1830)$ (6$\pm$2)%
$\frac{5}{2}^{+}$ $N(1680)$ (65$\pm$5)$\Sigma(1915)$ (10$\pm$5)$\Lambda(1820)$ (60$\pm$5)%
$\frac{7}{2}^{+}$ $\Delta(1950)$ (40$\pm$5)$\Sigma(2030)$ (20$\pm$3)%
The decays of octet and decuplet baryons into octet or decuplet baryon plus a pseudoscalar meson are related by SU(3). Such analyses have been performed by Samios, Goldberg, and Meadows in 1974 [594] and 30 years later by Guzey and Polyakov [595]. Both analyses agree that the data are compatible with SU(3) even though their assignments of baryons to multiplets are sometimes different and differ from our assignment as well.
One can also use arguments based on cascade decays of baryon resonances. Figure 3.9 shows a Dalitz plot for events due to the reaction $\gamma p\to p\pi^{0}\pi^{0}$ for the mass range $21001900<\sqrt{s}<$ 2100 MeV. Clear bands are visible. They are understood as decays of baryons with masses of about 2 GeV into intermediate resonances, e.g. a $N(1520)\nicefrac{{3}}{{2}}^{-}$, plus a pion; the intermediate resonance then decays into proton plus a second pion.
Refer to caption    Refer to caption Refer to caption
Figure 3.9: Left: $\gamma p\to p\pi^{0}\pi^{0}$ Dalitz plot for a selected $E_{\gamma}$ bin of 1975–2225 MeV (CBELSA/TAPS) [475]. The $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $N(1520)\nicefrac{{3}}{{2}}^{+}$, and $N(1680)\nicefrac{{5}}{{2}}^{+}$ are easily identified as intermediate states. Center: Cascade decays of resonances via an immediate state. Right: Classical orbits of nucleon excitations with $L=2$ (upper row) and $L=1$ (middle row) [596]. The first two pictures in the upper row show excitations of the $\rho$ and $\lambda$ oscillators, while in the third picture both $\rho$ and $\lambda$ are excited. When both oscillators are excited, de-excitation leads to an excited intermediate state (middle row).
In [110] it was shown that cascade decays via intermediate resonances with intrinsic orbital angular momentum stem preferentially from baryons with a contribution in the wave function of the excited state in which both oscillators are excited simultaneously. Resonances like, e.g., the $N(1900)\nicefrac{{3}}{{2}}^{+}$ have a mixed-symmetry spin-flavor wave function and both components $\phi_{00,20}-\phi_{00,02}$ and $\phi_{00,11}$ contribute to the spatial wave function. When a pion is emitted from the latter part from one oscillator, the other oscillator remains excited and de-excites in a second step. A coherent emission of the pion from both oscillators is, however, possible because then the two pions can interact and form a $\sigma$ meson. In resonances like the $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$, the $\rho$ and the $\lambda$ oscillators are excited coherently to $\phi_{00,20}+\phi_{00,02}$ which decays preferentially into two hadrons without orbital excitations. Both hypotheses hold true only approximately. The two oscillators can decay coherently, e.g. into $\pi N(1520)\nicefrac{{3}}{{2}}^{-}$, and this intermediate state can re-scatter into $\rho N$. On the other hand, resonances like the $\Delta(1950)\nicefrac{{7}}{{2}}^{-}$ can decay into $\pi N(1520)\nicefrac{{3}}{{2}}^{-}$ in a $D$ wave.

### v4

@@ 3.3.4   Beyond the third shell
3.3.5  Hyperons
The simplest spectrum is that of the $\Omega$ family. All $\Omega^{*}$’s belong to the decuplet, which has fewer resonances than octet baryons; they all have charge 1-1, and they are narrower: the $\Omega(2012)$ with unknown $J^{P}$ (but most likely $\nicefrac{{1}}{{2}}^{-}$) has a width of less than 10 MeV, compared to 110 MeV for the $\Delta(1620)\nicefrac{{1}}{{2}}^{-}$. However, experimentally very little is known: Only four excited states are known, and two of them likely belong to higher bands.
Table 3.7 presents a quark-model interpretation of the $\Lambda$ and $\Sigma$ resonances Klempt:2020bdu analogous to Fig. 3.2. In empty slots, no resonance is expected. Most exciting are the $\Lambda^{*}$’s, the only family that contributes to the flavor singlet spectrum. For each $N^{*}$ resonance, one $\Lambda^{*}$ octet state is expected. The $\Sigma^{*}$ resonances can belong to the $SU(3)_{f}$ octet or decuplet, so for each $N^{*}$ or $\Delta^{*}$ a $\Sigma^{*}$ resonance should exist. New data are expected from CLAS12 at JLab Burkert:2020akg , Ernst:2020hru PANDA at GSI/FAIR PANDA:2020hmi , and ELSA Cluster:2025 . For a recent review on hyperons we refer to Ref. Crede:2024hur .
Figure 3.7: (color online) Hyperon spectrum in the $SU(6)$ multiplet arrangement analogous to Fig. 3.2. In the first and second excitation band, all states expected from the quark model are shown, while the third band lists only multiplets for which at least one $\Lambda$ or $\Sigma$ candidate exists. The states with a dagger $\dagger$ are special: In the $\Lambda$ sector at about 2100 MeV, two states with $J^{P}=\nicefrac{{5}}{{2}}^{-}$ and $\nicefrac{{7}}{{2}}^{-}$ are expected with a mass separation of about 100 MeV but only one pair is known. In the $\Sigma$ sector, one pair is expected at about 1750 to 1800 MeV and two pairs at about 2000 to 2050 MeV. However, only two pairs are found and the third pair is missing. The observed pairs of states are possibly each mixtures of these configurations Klempt:2020bdu .
3.3.6  Baryon decays
Quark model calculations do not only give masses and widths of baryon resonances but also their wave functions (not only the harmonic oscillator basis wave functions). These can be used to calculate branching ratios, photocouplings or the $Q^{2}$ dependence of transition form factors. Table 3.3.6 gives a survey of references to calculations. A comprehensive review of the branching ratio calculations published before 2000 can be found in Capstick:2000qj .
In baryon decays $N^{*}\to NM$, a new $q\bar{q}$ pair is created in a process as depicted in Fig. 3.8. The antiquark combines with a quark of the resonance to form a meson. The newly created $q\bar{q}$ pair does not form the meson $M$, which is forbidden by the OZI rule Okubo:1963fa , Zweig:1964jf , Iizuka:1966fk . Often, the new $q\bar{q}$ pair is supposed to have the quantum numbers of the vacuum, $J^{PC}=0^{++}$. These $^{3}P_{0}$ models have met with considerable success.
Simple rules are at most approximately consistent with the experimental results. In hyperons, for example, the probability to emit a kaon is only 1/3, while the $s$ quark ends up in the baryon with a probability of 2/3. This makes the identification of hyperons in $K^{-}p$ elastic scattering difficult. In Table 3.3.6 we compare $N\bar{K}$ decays of $\Sigma$ hyperons with 4* rating with $N\pi$ decays of their non-strange partners. The suppression of $\Sigma^{*}\to N\bar{K}$ decays relative to their non-strange partners varies from 6 to 1.
In Refs. Fernandez-Ramirez:2015fbq , JPAC:2018zjz , a novel attempt is made to gain insight into the internal structure of baryons by modeling real and imaginary parts of baryon resonances. For the fits, baryons are divided into several classes: $N$, $\Delta$, $\Lambda$, and $\Sigma$ on the one hand, and according to their naturality and signature [^2] The naturality is given by $\eta=+1$ if $P=(-1)^{J_{p}-1/2}$ and $\eta=-1$ if $P=-(-1)^{J_{p}-1/2}$; the signature $\tau$ by $\tau=\nicefrac{{P}}{{\eta}}$. on the other hand. The real parts of resonances in a given class show the expected linearity as functions of $J$, the imaginary parts a square root behavior. The square-root dependence is assigned to the contribution of the phase space to the scattering amplitude, which is proportional to the momentum $q\sim\sqrt{s-s_{t}}$. The hope was that major deviations from the observed pattern would signal important components beyond compact $3q$ physics. The authors of Ref. Fernandez-Ramirez:2015fbq , JPAC:2018zjz conclude that in the $\Lambda$ and $\Sigma$ sector a consistent picture emerges for the leading hyperon Regge trajectories. The $\Lambda(1405)$ will be discussed below (see Section 6.2), where its two-pole nature is discussed. In Fernandez-Ramirez:2015fbq , it is found that the higher-mass pole of the $\Lambda(1405)$ fits into the general picture but the lower-mass pole does not, so it is suggested that it might be a molecule. The $N(1680)\nicefrac{{5}}{{2}}^{+}$, $N(1720)\nicefrac{{3}}{{2}}^{+}$, $\Delta(1930)\nicefrac{{3}}{{2}}^{-}$, and $\Delta(2400)\nicefrac{{9}}{{2}}^{-}$ are suggested to have sizable physics beyond the compact $3q$ picture. However, a few open questions remain: The $N$ and $\Delta$ are well-separated while the octet and decuplet $\Sigma$ resonances are not, and the decuplet $\Sigma(1385)\nicefrac{{3}}{{2}}^{+}$ is shown jointly with the octet $\Sigma(1775)\nicefrac{{5}}{{2}}^{-}$ (these quantum numbers are forbidden for a decuplet state) and with the decuplet $\Sigma(2030)\nicefrac{{7}}{{2}}^{+}$. The $N(1720)\nicefrac{{3}}{{2}}^{+}$ is related to the $N(1990)\nicefrac{{7}}{{2}}^{+}$; both belong to the second excitation shell, and both have $L=2$ as their leading orbital angular momentum. The $\Delta(1930)\nicefrac{{3}}{{2}}^{-}$ must have one unit of radial excitation, and the $\Delta(2400)\nicefrac{{9}}{{2}}^{-}$ likely has the same. It seems questionable whether these states should be related to the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$, and $\Delta(2420)\nicefrac{{11}}{{2}}^{+}$.
$J^{P}$ $B^{*}\to N\pi$ $B^{*}\to N\bar{K}$ $B^{*}\to N\bar{K}$
[…]
The decays of octet and decuplet baryons into octet or decuplet baryon plus a pseudoscalar meson are related by SU(3). Such analyses have been performed by Samios, Goldberg, and Meadows in 1974 Samios:1974tw and 30 years later by Guzey and Polyakov Guzey:2005vz . Both analyses agree that the data are compatible with SU(3) even though their assignments of baryons to multiplets are sometimes different and differ from our assignment as well.
One can also use arguments based on cascade decays of baryon resonances. Figure 3.9 shows a Dalitz plot for events due to the reaction $\gamma p\to p\pi^{0}\pi^{0}$ for the mass range $1900<\sqrt{s}<2100$ MeV. Clear bands are visible. They are understood as decays of baryons with masses of about 2 GeV into intermediate resonances, e.g. a $N(1520)\nicefrac{{3}}{{2}}^{-}$, plus a pion; the intermediate resonance then decays into proton plus a second pion.
Refer to caption   Refer to caption Refer to caption
Figure 3.9: Left: $\gamma p\to p\pi^{0}\pi^{0}$ Dalitz plot for a selected $E_{\gamma}$ bin of 1975–2225 MeV (CBELSA/TAPS) Sarantsev:2025lik . The $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $N(1520)\nicefrac{{3}}{{2}}^{+}$, and $N(1680)\nicefrac{{5}}{{2}}^{+}$ are easily identified as intermediate states. Center: Cascade decays of resonances via an immediate state. Right: Classical orbits of nucleon excitations with $L=2$ (upper row) and $L=1$ (middle row) CBELSATAPS:2015taz . The first two pictures in the upper row show excitations of the $\rho$ and $\lambda$ oscillators, while in the third picture both $\rho$ and $\lambda$ are excited. When both oscillators are excited, de-excitation leads to an excited intermediate state (middle row).
In Thiel:2013cea it was shown that cascade decays via intermediate resonances with intrinsic orbital angular momentum stem preferentially from baryons with a contribution in the wave function of the excited state in which both oscillators are excited simultaneously. Resonances like, e.g., the $N(1900)\nicefrac{{3}}{{2}}^{+}$ have a mixed-symmetry spin-flavor wave function and both components $\phi_{00,20}-\phi_{00,02}$ and $\phi_{00,11}$ contribute to the spatial wave function. When a pion is emitted from the latter part from one oscillator, the other oscillator remains excited and de-excites in a second step. A coherent emission of the pion from both oscillators is, however, possible because then the two pions can interact and form a $\sigma$ meson. In resonances like the $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$, the $\rho$ and the $\lambda$ oscillators are excited coherently to $\phi_{00,20}+\phi_{00,02}$ which decays preferentially into two hadrons without orbital excitations. Both hypotheses hold true only approximately. The two oscillators can decay coherently, e.g. into $\pi N(1520)\nicefrac{{3}}{{2}}^{-}$, and this intermediate state can re-scatter into $\rho N$. On the other hand, resonances like the $\Delta(1950)\nicefrac{{7}}{{2}}^{-}$ can decay into $\pi N(1520)\nicefrac{{3}}{{2}}^{-}$ in a $D$ wave.
@@ 3.3.7  Comments
[…]
@@
The good agreement between the non-relativistic quark model predictions and the measured masses is, and has always been, extremely astonishing. Even constituent quarks should move at relativistic velocities when the nucleon is excited. Furthermore, the coupling of baryons to virtual baryon-meson intermediate states may lead to large baryon self-energies and to mixing of baryons of the same quantum numbers Morel:2002vk . There is no reason at all to believe that the non-relativistic quark model may be useful in understanding the spectrum, except one: it works surprisingly well. Possibly, the concept of rapidly moving constituent quarks with a fixed rest mass is misleading: when a gluon string evolves, a string of confined color field connecting quarks, a large fraction of the excitation energy may be stored in the string, the effective constituent quark masses increase, and the resulting motion could remain a classical one.
@@
In Ref. Anisovich:1998au , two resonances were considered with one decay channel only, in which case one of the states becomes narrow and the other one broad. We know, however, no physical case corresponding to this situation. The second possibility are two close-by states with strong couplings to two decay modes. For example, the two mesons $K_{1A}$ and $K_{1B}$, the strange partners of $a_{1}(1260)$ and $b_{1}(1235)$, have the decay modes $K^{*}\pi$ and $K\rho$ which are expected with similar rates. Due to mixing, these two mesons repel each other, and the mixed state $K_{1}(1280)$ couples more strongly to $K\rho$ and $K_{1}(1400)$ very strongly to $K^{*}\pi$. On the other hand, these two scenarios – decays into one dominant channel or mutual repulsion of pairs of states – do not seem to be applicable in the high-lying baryon spectrum, where the number of negative-parity $N^{*}$ and $\Delta^{*}$ states expected in the third excitation shell is rather large (see Table 3.2). In the $N^{*}$ sector there is a low-mass doublet and in the $\Delta^{*}$ sector a low-mass triplet, and only these five low-mass resonances can be assigned to a unique multiplet. Above, there is exactly one observed state per partial wave from $J^{P}=\nicefrac{{1}}{{2}}^{-}$ to $\nicefrac{{9}}{{2}}^{-}$. In the quark model, the four $\Delta\nicefrac{{3}}{{2}}^{-}$ states above the $\Delta(1940)\nicefrac{{3}}{{2}}^{-}$ in the third excitation shell have calculated masses within a 100 MeV wide interval and the seven $N\nicefrac{{3}}{{2}}^{-}$ states above the $N(1875)\nicefrac{{3}}{{2}}^{-}$ fall into a 200 MeV wide interval, and these resonances are expected to have many decay modes.